OpenStudy (amonoconnor):

I've worked through this problem several times, but can't figure out how to get the answer the book has, or what I'm doing wrong. Can someone take a look at my process, and explain how to do it "right"? What am I messing up or forgetting? What should I be using? The problem: (x^4)(x+y) = (y^2)(3x-y); Find dy/dx using Implicit Differentiation.

OpenStudy (amonoconnor):

This is how I've gone about finding dy/dx:

OpenStudy (amonoconnor):

\[x^4(x+y) = y^2(3-y) \] \[(4x^3)(x+y)+x^4(1+(dy/dx)=2y(dy/dx)(3x-y) + y^2(3(1)-(dy/dx))\] \[4x^4 + 4x^3y+x^4+x^4(dy/dx) = 6xy(dy/dx) -2y^2(dy/dx)+3y^2-y^2(dy/dx)\] \[5x^4+4x^3y+x^4(dy/dx) = 6xy(dy/dx) +3y^2 - 3y^2(dy/dx)\]

OpenStudy (amonoconnor):

\[5x^4+4x^3y+x^4(dy/dx)=6xy(dy/dx)-3y^2(dy/dx)+3y^2\] \[5x^4+4x^3y+-3y^2=6xy(dy/dx)-3y^2(dy/dx)-x^4(dy/dx)\] \[5x^4+4x^3y-3y^2=dy/dx(6xy-3y^2-x^4)\] \[dy/dx = \frac{5x^4+4x^3y-3y^2}{6xy-3y^2-x^4}\]

OpenStudy (amonoconnor):

*Sorry for the last step looking funky – OS has been making fractions look off for a couple of weeks now at least... ?

OpenStudy (johnweldon1993):

No that's fine.. And everything looks perfect to me...what does the book have?

OpenStudy (amonoconnor):

The book has: \[\frac{3y^2-5x^4-4x^3y}{x^4+3y^2-6xy}\]

OpenStudy (amonoconnor):

And I initially was like "Oh, I just multiply the top and bottom by -1, right?" But I must ask... why would I do that, and wouldn't that make it a NEGATIVE dy/dx then?

OpenStudy (freckles):

-1/-1=1

OpenStudy (johnweldon1993):

Yes thank you @freckles lol I couldnt type it quick enough

OpenStudy (amonoconnor):

Exactly, so I'm like, super confused about how the signs are all switched, but the terms are the same...

OpenStudy (freckles):

Your answer is fine though. They probably did there adding and subtracting of things on both sides a bit differently than you. But yeah the answers are equal.

OpenStudy (amonoconnor):

You guys sure? That was my gut reaction too when I read the book, but I don't want to get no credit, if it's some technical thing.

OpenStudy (freckles):

Yes I'm sure - 1/-1 is 1

OpenStudy (johnweldon1993):

Do it...multiply BOTH top and bottom by -1 it will change all the signs you want And you're overthinking it...it will NOT equal -dy/dx *Because as @freckles is saying...-1/-1 = 1

OpenStudy (johnweldon1993):

-dy/dx would be just changing the signs on either top or bottom...not both

OpenStudy (amonoconnor):

But you're sure that both final answers are mathematically equivalent, that's what I'm pondering the truth of... ? Alright. But why would it make sense to do this in this case? Where does it come from, if I worked my math out fine? If I had never looked at the back of the book, I would have assumed I had the right answer and would have done everything I could have, ostensibly being "done", right? How would I know to change the signs like this?

OpenStudy (johnweldon1993):

You wouldn't need to, they are mathematically equivalent...Admittedly, this does always throw me off as well *when a textbook has SUCH a similar answer in a different manner* however, in this case, they are the same

OpenStudy (amonoconnor):

Thank you both, very much, I really appreciate the clarification :) It helps me learn sooo much better

OpenStudy (johnweldon1993):

No problem!

OpenStudy (freckles):

here is an example of why you are equivalently asking if -1/-1=1 say we have the equation: -x+5=-2+2x add x and add 2 on both sides giving us: 5+2=x+2x 5+2=x(1+2) so x=(5+2)/(1+2) or just 7/3 now if I chose to subtract 2x and subtract 5 on both sides that would give us: -x-2x=-2-5 x(-1-2)=-2-5 x=(-2-5)/(-1-2) which is the just -7/-3 which is just 7/3 So I was saying earlier they probably didn't take their answer and multiply -1/-1. They probably just went about solving for dy/dx differently by choosing to add/subtract different terms than you on both sides.